OpenStudy (fionacndg):

Find the area of the shaded region. Use 3.14 for , and round your answer to the nearest tenth.

OpenStudy (midhun.madhu1987):

figure?

OpenStudy (fionacndg):

@midhun.madhu1987

OpenStudy (midhun.madhu1987):

Area of Rectangle = length x breadth = 6 x 5 = 30 Area of Semi Circle = 1/2(pi*r*r) = 0.5 * 3.14 * 2.5 * 2.5 = 9.8 Area of remaining part = area of rectangle - area of semi circle = 30 - 9.8 = 20.2

OpenStudy (fionacndg):

so here are the answers, a.20.2 b.10.4 c.25.5 d.13.3 would A be the answer then? @midhun.madhu1987

OpenStudy (midhun.madhu1987):

yes

OpenStudy (fionacndg):

thank you!